Calculate the number of mole presents in certain mass of gas occupying 6.5dm3 at atm and 15•C. ( R=0.082 atm dm3 K-1 mole-1).
1
Expert's answer
2020-01-08T02:02:14-0500

Task:

Calculate the number of mole presents in certain mass of gas occupying 6.5 dm3 at ??? atm and 15C. (R = 0.082 atm * dm3 * K-1 * mol-1).


Solution:

The ideal gas equation is:

pV = nRT


p = ??? atm = X atm

V = 6.5 dm3;

R = 0.082 atm * dm3 * K-1 * mol-1;

T = 15C + 273 = 288 К.

Then,

n = pV / RT = (X * 6.5) / (0.082 * 288) = 6.5X / 23.616 = 0.2752*X moles.

For example, if p = 1 atm, then: n = 0.2752 * 1 = 0.2752 moles.


Answer: 0.2752*X moles of gas.